# Self-Tuning: Instructing LLMs to Effectively Acquire New Knowledge through Self-Teaching

### v3

@@ Appendix D Qualitative Analysis
In Table D, we provide a qualitative comparison between the answers generated by PIT and Self-Tuning on the Wiki-Bio test set. We observe that Self-Tuning performs better in answering questions that inquire about multiple facts and time-related facts, as indicated in the top part of Table D. Furthermore, as shown in the lower part, PIT tend to fail to recall and extract facts at the end of the documents, i.e., suffering from “positional bias”. This observation is consistent with the findings in Saito et al. (2024). Encouragingly, our proposed Self-Tuning aids in recalling and extracting factual knowledge embedded at the end of the documents. These findings align with the automatic evaluation results, underscoring the effectiveness of Self-Tuning in enhancing the LLM’s knowledge acquisition capability, particularly in knowledge extraction.
Appendix E Error Analysis
@@
Appendix F In-depth Sample Documents and Corresponding QA Pairs for Open-Ended Generation and Natural Language Inference Tasks
@@
Appendix G Token Count Distribution for the Open-ended Generation Task Across the Three Datasets
@@
Appendix H Examination of QA Types in Open-ended Generation QA Datasets
@@
We perform a detailed analysis of the QA types associated with the factual information in the open-ended generation QA datasets, as displayed in Table H, by using the prompt in Table 21 with GPT-4.
Appendix I Detailed Templates used in the Self-Teaching Strategy
@@
Appendix J Datasets and Evaluation Metrics
@@
We assess the effectiveness of Self-Tuning in enhancing the model’s knowledge acquisition capabilities on the curated Wiki-Newpages-QA datasets, concentrating on memorization, extraction, and reasoning. For memorization, we utilize test document datasets and report perplexity Jelinek et al. (1977), which measures how well a language model predicts a text sample. For extraction, we employ test QA datasets for open-ended generation tasks. To evaluate the factual accuracy of the generated responses, we use exact match (EM), Recall, and F1 over words in the answer(s), following Kwiatkowski et al. (2019). Additionally, we report Rouge-L Lin (2004) to measure the overlap of n-grams between the generated and gold answers, accounting for minor lexical variations, following Jiang et al. (2024b). We also assess accuracy by comparing each response’s factual correctness to the gold answer, using the bidirectional entailment approach with the Deberta-Large-MNLI model He et al. (2021). We report the five-shot evaluation results on the open-ended generation tasks using the prompt in Table 20. Concerning reasoning, we utilize the test QA datasets for NLI tasks and report the accuracy by comparing the generated option with the gold option using EM. We present the zero-shot evaluation results on NLI tasks.
@@
It is well-known that knowledge acquisition is often accompanied by catastrophic forgetting Allen-Zhu and Li (2023); Wang et al. (2023). Therefore, we also provide the knowledge retention performance for a comprehensive investigation. Specifically, we verify the knowledge extraction performance on world knowledge using natural questions (NQ) Kwiatkowski et al. (2019) (i.e., NQ-open Min et al. (2021) in the closed-book setting) and report EM and F1 scores. We report the five-shot evaluation results using the first five QA pairs in the dev sets as prompts. we assess the reasoning capability on Commonsense knowledge using CommonsenseQA (CSQA) Talmor et al. (2019), employing accuracy to assess the correctness of the selected option, calculated by comparing the generated option against the gold option using EM. We present the five-shot performance on the dev sets, as the test set does not contain golden annotations, and use the first five multi-choice QA pairs in the training set as prompts. We use these two datasets because they were curated before the cut-off time of Llama2 family models (i.e., year 2022), making it likely that the models have obtained relevant knowledge in these datasets during the pre-training stage, as evidenced by Touvron et al. (2023).
Appendix K Implementation Details
@@
We train Llama2-7B and Llama2-7B-chat on 8 32GB Tesla V100 GPUs using a batch size of 8 and a learning rate of 5e-6. Additionally, we train Llama2-13B on 8 A100-SXM4-40GB GPUs with a batch size of 8 and a learning rate of 5e-6. To ensure a fair comparison, all compared approaches train on the test documents for 3 epochs in total, regardless of the number of training stages. For continued pre-training, which is observed to struggle in grasping new knowledge, we train the models for 5 epochs. The specific number of training epochs used for each approach in Table 8 are as follows:
@@
PIT Jiang et al. (2024b) first trains on and datasets for 3 epochs, positioning the QA pairs right before the corresponding document texts, then trains on the data for 3 epochs.
@@
Specifically, in the cross-domain setting, where there is a substantial difference between the domains of the training data and test documents, we continue training on data while reviewing the data for 2 epochs after the initial training stage, followed by further training on data for 1 epoch. Furthermore, we adopt the same training strategy when dealing with Llama2-7B-chat, where the process of knowledge injection poses a significant challenge, as demonstrated by our experimental results. In accordance with Jiang et al. (2024b), for PIT and Self-Tuning, we include 64 examples and 128 examples randomly sampled from datasets, respectively, during the final training stages when solely training on the data, to prevent the model from losing its question-answering capabilities. It is important to note that all evaluation results are reported at the temperature .
@@
Standard Instruction-Tuning w/o Forgetting Jiang et al. (2024b) initially trains on the mixture of and for 3 epochs, then on and datasets for 1 epoch.
PIT++ Jiang et al. (2024b) initially trains on for 1 epoch, then on and datasets for 3 epochs, with the QA pairs placed right before the corresponding document texts, and finally, it trains on the data for 3 epochs.
Mixed Training Jiang et al. (2024b) trains on mixture of the , and datasets simultaneously for 3 epochs.
@@
Appendix L A Sample Training Document in the Reading-Comprehension Format

### v4

@@ Appendix E Qualitative Analysis
In Table E, we provide a qualitative comparison between the answers generated by PIT and Self-Tuning on the Wiki-Bio test set. We observe that Self-Tuning performs better in answering questions that inquire about multiple facts and time-related facts, as indicated in the top part of Table E. Furthermore, as shown in the lower part, PIT tend to fail to recall and extract facts at the end of the documents, i.e., suffering from “positional bias”. This observation is consistent with the findings in Saito et al. (2024). Encouragingly, our proposed Self-Tuning aids in recalling and extracting factual knowledge embedded at the end of the documents. These findings align with the automatic evaluation results, underscoring the effectiveness of Self-Tuning in enhancing the LLM’s knowledge acquisition capability, particularly in knowledge extraction.
Appendix F Ablation Study
Appendix G Error Analysis
@@
Appendix H Evaluation Results on Self-Tuning Variants
| Mem. | Extraction | Reason. | Extraction | Reasoning | ||||||
[…]
Setup. To further investigate the effectiveness of Self-Tuning, we present three variants, as depicted in Table 9: (1) Self-Tuning w/o Review, where we continue training on test documents without the reviewing capability; (2) Self-Tuning via Read., which displays the training documents in a reading-comprehension format Cheng et al. (2024) (an example is shown in Table 28); (3) Self-Tuning w/ Pre-Review, which trains on a combination of training documents and training QA in the second stage, before training on test documents.
Results. In Table H, despite having lower performance than Self-Tuning, all variations significantly enhance the model’s ability for knowledge acquisition compared to continued pre-training, which further validates the effectiveness of Self-Tuning in improving knowledge acquisition.
Reviewing QA ability aids in both knowledge acquisition and retention. Compared to Self-Tuning, Self-Tuning w/o Review also displays inferior performance on the knowledge retention task.
Appendix I Evaluation Results on Llama2-13B in the Single-domain Scenario
| Memorization | Extraction | Reason. | Extraction | Reasoning | ||||||
| Llama2-13B | ||||||||||
| w/o Knowledge Injection | ||||||||||
| w/ Knowledge Injection | ||||||||||
Table 13 presents the evaluation results on Llama2-13B concerning knowledge acquisition and retention in the single-domain scenario using the Wiki-Bio dataset. We make the following observations:
Self-Tuning consistently demonstrates superior performance in enhancing the model’s knowledge acquisition and retention abilities as the model size scales. As the model size scales, Self-Tuning continues to achieve the highest performance across all evaluation metrics on memorization and acquisition tasks, consistently outperforming the compared methods by a significant margin (e.g., improving EM score by 20% on the extraction task). On the reasoning task, Self-Tuning consistently attains high accuracy. Additionally, Self-Tuning consistently exhibits strong performance on knowledge retention tasks. These findings confirm the effectiveness of Self-Tuning, suggesting the potential and robustness of Self-Tuning for applications on larger-scale models.
Continued pre-training for knowledge acquisition proves challenging across all three dimensions. We find that continuing pre-training on new documents may result in a decline in knowledge extraction performance on Llama2-13B, compared to the baseline performance. This could be due to the fact that merely continuing pre-training might adversely affect its question-answering capability, even when equipped with new knowledge, as demonstrated by the lowered PPL. This observation is consistent with the findings in Cheng et al. (2024). Moreover, the marginal improvements in memorization (reducing PPL by 2%) and reasoning (increasing accuracy by 2%) suggest that such a naive approach fails to help the model memorize and capitalize on new knowledge. This highlights the importance of evaluating the model’s knowledge acquisition ability comprehensively across multiple dimensions.
Appendix J Evaluation Results on Llama2-7B-chat in the Single-domain Scenario
| Memorization | Extraction | Reason. | Extraction | Reasoning | ||||||
| Llama2-7B-chat | ||||||||||
| w/o Knowledge Injection | ||||||||||
| w/ Knowledge Injection | ||||||||||
Appendix K Training Efficiency Analysis
To ensure a fair comparison, all methods for knowledge injection presented in Table 3 were trained on raw test documents for 3 epochs, as detailed in Appendix S. Additionally, we conducted a detailed analysis of training efficiency on the Wiki-Newpages-2023-Bio dataset using 8 Tesla V100 GPUs (32G) with Llama2-7B:
Continued pre-training: 112.91 seconds
Standard instruction-tuning: 1661.06 seconds
PIT: 6205.52 seconds
Self-Tuning: 5220.50 seconds
Our Self-Tuning significantly outperforms the most competitive baseline, PIT, on the knowledge acquisition task and is more time-efficient.
Appendix L Evaluation Results on Gemma-7B in the Single-Domain Scenario
| Method | Knowledge Acquisition | |||||
| w/o Knowledge Injection | ||||||
| Closed-book | 12.41 | 7.09 | 4.68 | 17.60 | 18.10 | 17.65 |
| w/ Knowledge Injection | ||||||
| Continued Pre-training | 3.99 | 8.14 | 6.33 | 19.91 | 20.97 | 19.82 |
| Standard Instruction-tuning | 10.13 | 10.41 | 8.60 | 24.06 | 23.86 | 24.16 |
| PIT | 4.19 | 8.14 | 5.88 | 20.87 | 20.78 | 20.68 |
| Self-Tuning | 1.09 | 41.93 | 36.80 | 56.95 | 57.41 | 56.30 |
We present the evaluation results for Gemma-7B in Table 15. Our Self-Tuning method consistently achieves the best performance, significantly outperforming the baseline methods by a substantial margin. This observation aligns with the results reported across all other evaluation scenarios in Section 5.
Appendix M Evaluation Results with Continual Learning Techniques
| Method | Wiki-Newpages-2023-QA (Acquisition) | NQ (Reten.) | CSQA (Reten.) | |||||||
| Memorization | Extraction | Reason. | Extraction | Reasoning | ||||||
| Knowledge Acquisition on Wiki-Newpages-2023-10-Bio (Single-Domain Scenario) | ||||||||||
| w/o Knowledge Injection | ||||||||||
| w/ Knowledge Injection | ||||||||||
| Self-Tuning+Replay | 1.03 | 44.49 | 39.82 | 58.44 | 60.58 | 58.00 | 56.24 | 22.67 | 33.86 | 73.55 |
This section explores the potential of integrating Self-Tuning with continual learning techniques. Table 16 presents the evaluation results of combining Self-Tuning with a representative continual learning strategy: the replay-based method. In this approach, 500 training QA pairs were randomly sampled from the Wiki QA datasets, curated prior to the pre-training cutoff date of Llama2-7B Zhang et al. (2024c). These QA pairs were included throughout the training process to reinforce general domain knowledge.
The evaluation results on MCQA and NQ confirm that Self-Tuning effectively preserves previously acquired knowledge. Moreover, integrating Self-Tuning with replay-based continual learning (Self-Tuning+Replay) further enhances model performance, demonstrating the following benefits:
1.
Effective knowledge acquisition: The model successfully learns new knowledge from previously unseen raw documents. This aligns with findings in Appendix H, which highlight that the reviewing QA ability facilitates knowledge acquisition.
2.
Efficient knowledge retention: The replay-based approach ensures that previously learned knowledge is preserved, mitigating catastrophic forgetting during the learning of new tasks.
These findings underscore the significant potential of integrating Self-Tuning with continual learning techniques. By combining Self-Tuning’s strengths with replay-based strategies, the model not only excels in acquiring new knowledge but also maintains strong retention of existing information, making this approach an effective solution for long-term knowledge management.
Appendix N Evaluation Results Comparing with a Baseline Utilizing Document-Based QA Generation on Test Corpora
| Method | Knowledge Acquisition (Wiki-Newpages-2023-Bio) | |||||
| w/o Knowledge Injection | ||||||
| Open-book w/ test doc | 8.41 | 55.20 | 31.83 | 64.48 | 75.55 | 62.10 |
| Closed-book | 8.41 | 4.68 | 2.87 | 14.63 | 16.98 | 15.07 |
| w/ Knowledge Injection | ||||||
| Continued Pre-training | 7.28 | 6.33 | 3.62 | 15.96 | 18.72 | 16.11 |
| Training on test doc w/ QA pairs | 1.08 | 15.84 | 12.07 | 28.58 | 31.06 | 28.07 |
| Self-Tuning | 1.11 | 37.25 | 31.52 | 50.83 | 52.62 | 50.61 |
Constructing QA data for every newly introduced raw corpus is infeasible in real-world scenarios due to cost and scalability constraints. Our goal is to enable the model to autonomously acquire new knowledge from previously unseen raw test documents. Therefore, approaches that rely on constructing and training on QA pairs specifically tailored to test documents, as explored in prior studies Mecklenburg et al. (2024); Jiang et al. (2024a), are not aligned with the objectives of this work.
Nevertheless, for completeness and to provide additional context, we include the results of a baseline that simultaneously trains on test documents and QA pairs generated using our proposed Self-Teaching strategy.
As presented in Table 17, Self-Tuning consistently demonstrates superior performance. This comparison highlights the significant advantages of our approach in fostering autonomous knowledge acquisition without depending on pre-constructed QA pairs tailored to the test data.
@@ Appendix I Evaluation Results on Llama2-13B in the Single-domain Scenario
Appendix O In-depth Sample Documents and Corresponding QA Pairs for Open-Ended Generation and Natural Language Inference Tasks
@@
Appendix P Token Count Distribution for the Open-ended Generation Task Across the Three Datasets
@@
Appendix Q Examination of QA Types in Open-ended Generation QA Datasets
@@
We perform a detailed analysis of the QA types associated with the factual information in the open-ended generation QA datasets, as displayed in Table Q, by using the prompt in Table 26 with GPT-4.
Appendix R Detailed Templates used in the Self-Teaching Strategy
@@
Appendix S Datasets and Evaluation Metrics
@@
We assess the effectiveness of Self-Tuning in enhancing the model’s knowledge acquisition capabilities on the curated Wiki-Newpages-QA datasets, concentrating on memorization, extraction, and reasoning. For memorization, we utilize test document datasets and report perplexity Jelinek et al. (1977), which measures how well a language model predicts a text sample. For extraction, we employ test QA datasets for open-ended generation tasks. To evaluate the factual accuracy of the generated responses, we use exact match (EM), Recall, and F1 over words in the answer(s), following Kwiatkowski et al. (2019). Additionally, we report Rouge-L Lin (2004) to measure the overlap of n-grams between the generated and gold answers, accounting for minor lexical variations, following Jiang et al. (2024c). We also assess accuracy by comparing each response’s factual correctness to the gold answer, using the bidirectional entailment approach with the Deberta-Large-MNLI model He et al. (2021). We report the five-shot evaluation results on the open-ended generation tasks using the prompt in Table 25. Concerning reasoning, we utilize the test QA datasets for NLI tasks and report the accuracy by comparing the generated option with the gold option using EM. We present the zero-shot evaluation results on NLI tasks.
@@
It is well-known that knowledge acquisition is often accompanied by catastrophic forgetting Allen-Zhu and Li (2023); Wang et al. (2023). Therefore, we also provide the knowledge retention performance for a comprehensive investigation. Specifically, we verify the knowledge extraction performance on world knowledge using natural questions (NQ) Kwiatkowski et al. (2019) (i.e., NQ-open Min et al. (2021) in the closed-book setting) and report EM and F1 scores. We report the five-shot evaluation results using the first five QA pairs in the dev sets as prompts. we assess the reasoning capability on Commonsense knowledge using CommonsenseQA (CSQA) Talmor et al. (2019), employing accuracy to assess the correctness of the selected option, calculated by comparing the generated option against the gold option using EM. We present the five-shot performance on the dev sets, as the test set does not contain golden annotations, and use the first five multi-choice QA pairs in the training set as prompts. We use these two datasets because they were curated before the cut-off time of Llama2 family models (i.e., year 2022), making it likely that the models have obtained relevant knowledge in these datasets during the pre-training stage, as evidenced by Touvron et al. (2023a).
Appendix T Implementation Details
@@
We utilize Llama2-7B for our investigation and provide analyses on Qwen2-7B, Mistral-7B-v0.1, Gemma-7B, Llama2-13B, and Llama2-7B-chat for a comprehensive understanding. We use the following training objectives: for training on document data , we compute the standard next-token prediction loss by averaging over all tokens in the document (Equation 4); for training on QA data , we compute the average negative log-likelihood loss only on tokens in the answer given the question (Equation 5), where and refer to the length of the tokenized document sequence and answer sequence, respectively.
We train Llama2-7B, Qwen2-7B, Mistral-7B-v0.1, Gemma-7B, and Llama2-7B-chat on 8 32GB Tesla V100 GPUs using a batch size of 8 and a learning rate of 5e-6. Additionally, we train Llama2-13B on 8 A100-SXM4-40GB GPUs with a batch size of 8 and a learning rate of 5e-6. To ensure a fair comparison, all compared approaches train on the test documents for 3 epochs in total, regardless of the number of training stages. For continued pre-training, which is observed to struggle in grasping new knowledge, we train the models for 5 epochs. The specific number of training epochs used for each approach in Table 9 are as follows:
@@
PIT Jiang et al. (2024c) first trains on and datasets for 3 epochs, positioning the QA pairs right before the corresponding document texts, then trains on the data for 3 epochs.
@@
Specifically, in the cross-domain setting, where there is a substantial difference between the domains of the training data and test documents, we continue training on data while reviewing the data for 2 epochs after the initial training stage, followed by further training on data for 1 epoch. Furthermore, we adopt the same training strategy when dealing with Llama2-7B-chat, where the process of knowledge injection poses a significant challenge, as demonstrated by our experimental results. In accordance with Jiang et al. (2024c), for PIT and Self-Tuning, we include 64 examples and 128 examples randomly sampled from datasets, respectively, during the final training stages when solely training on the data, to prevent the model from losing its question-answering capabilities. It is important to note that all evaluation results are reported at the temperature .
@@
Standard Instruction-Tuning w/o Forgetting initially trains on the mixture of and for 3 epochs, then on and datasets for 1 epoch.
PIT++ Jiang et al. (2024c) initially trains on for 1 epoch, then on and datasets for 3 epochs, with the QA pairs placed right before the corresponding document texts, and finally, it trains on the data for 3 epochs.
Mixed Training trains on mixture of the , and datasets simultaneously for 3 epochs.
Future research could explore the inclusion of segments from general domain datasets, such as Wiki data Zhang et al. (2024c) and the Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2021), which were compiled prior to the pre-training cut-off date. Adopting this strategy may improve the model’s capacity to retain learned knowledge and skills while reducing the risk of overfitting to novel information. In our current study, we deliberately avoid integrating extra data to ensure a precise assessment of knowledge injection, thereby preventing any biases that might arise from the inclusion of additional sources.
@@
| The prompt utilized by GPT-4o for building QA datasets for open-ended generation tasks |
| Your task is to rephrase the paragraph below to make it clearer and more concise. Then, create a detailed list of questions and corresponding answers that cover the factual information in the revised content. Answers should be concise, consisting of a few short phrases separated by commas. For example: |
| Paragraph: |
| 6 VCs explain how startups can capture and defend marketshare in the AI era. Ninety-four percent of business leaders agree AI will be critical to all businesses’ success over the next five years, and total global spending on AI is expected to reach $154 billion by the end of this year, a 27% increase from 2022. |
| Revised Content: |
| Six venture capitalists (VCs) explain how startups can capture and defend market share in the AI era. Ninety-four percent of business leaders agree that AI will be critical to the success of all businesses over the next five years. Additionally, total global spending on AI is expected to reach $154 billion by the end of this year, representing a 27% increase from 2022. |
| Simple Question-Answering Pairs: |
| Question: How many VCs explain how startups can capture and defend market share in the AI era? |
| Answer: Six venture capitalists (VCs). |
| Question: What percentage of business leaders agree that AI will be critical to the success of all businesses over the next five years? |
| Answer: Ninety-four percent. |
| Question: Over what period do business leaders believe AI will be critical to all businesses’ success? |
| Answer: Over the next five years. |
| Question: How much is the total global spending on AI expected to reach by the end of this year? |
| Answer: $154 billion. |
| Question: By what percentage is the global spending on AI expected to increase from 2022? |
| Answer: Twenty-seven percent. |
| Your task is to rephrase the paragraph below to make it clearer and more concise. Then, create a detailed list of simple questions and corresponding answers that cover the information in the revised content. Answers should be concise, consisting of a few short phrases separated by commas. For example: |
| Paragraph: |
{paragraph}
| Revised Content: |
Appendix U Implementation Details of Evaluation on Varying Models and Corpora
U.1 Evaluation on Different Models
For the evaluation using different models, specifically Qwen2-7B Yang et al. (2024), we collected articles published on Wikipedia NewPages from June 2024 to September 2024 to minimize overlap with the pre-training corpus. We randomly selected 146 biographies from the collected articles, following the data construction pipeline described in Appendix B, to create a new question-answering dataset for an open-ended generation task. This resulted in a test set, named WikiBio-2024, comprising 146 documents and a total of 827 QA pairs.
U.2 Evaluation on Varied Corpora
For the evaluation using varied corpora, we utilized the news data collected by Tang and Yang (2024) using the mediastack API101010https://mediastack.com/. Specifically, this dataset includes articles published from September 26, 2023, to December 26, 2023, which is beyond the pre-training cutoff time of Llama2-7B. The dataset covers a range of news categories, such as entertainment, business, technology, and science.
For each factual sentence extracted from the original articles by Tang and Yang (2024), we concatenated the article title with the fact to create a knowledge snippet. Following Tang and Yang (2024), we used GPT-4o OpenAI (2024) (version dated 2024-02-01) to first paraphrase these snippets to make them clearer and more concise, and then generate relevant QA pairs. The prompt utilized can be found in Table 27. Using the data construction pipeline described in Appendix B, we generated a new training set, i.e., WebNews-2023, consisting of 1,800 training documents and 6,038 QA pairs, as well as a testing set with 400 testing documents and 1,350 QA pairs.
Appendix V A Sample Training Document in the Reading-Comprehension Format